| -3*4 - 12*2 | -8(-7)
Can you please simplify this absolute value in easy to understand steps.

Answer by richard1234(7193)

Do the absolute value part first.

The expression -3*4 - 12*2 is equal to -12 - 24, or -36. The absolute value of -36 is 36.

Now it's just doing 36 - 8(-7), which is equal to 36 - (-56) = 36 + 56 = 92.